Define solubility product constant and derive the relationships between solubility
and solubility product constants for salts of AB2, A2B types.
1
Expert's answer
2020-05-20T13:44:43-0400

Solution.

1.

"AB2 = A^{2+} + 2B^-"

"C(AB2) = Solubility = X"

"[A^{2+}] = X"

"[B^-] = 2 \times X"

"Ks = [A^{2+}] \times [B^-]^2 = X \times (2\times X)^2 = 4 \times X^3"

2.

"A2B = 2A^+ + B^{2-}"

"C(A2B) = Solubility = X"

"[A^+] = 2 \times X"

"[B^{2-}] = X"

"Ks = [A^+]^2 \times [B^{2-}] = (2 \times X)^2 \times X = 4 \times X^3"

Answer:

1.

"[A^{2+}] = X"

"[B^-] = 2 \times X"

"Ks = [A^{2+}] \times [B^-]^2 = X \times (2\times X)^2 = 4 \times X^3"

2.

"[A^+] = 2 \times X"

"[B^{2-}] = X"

"Ks = [A^+]^2 \times [B^{2-}] = (2 \times X)^2 \times X = 4 \times X^3"